OpenStudy (anonymous):

Simplify (4^(n+1))-(4^n)

OpenStudy (anonymous):

It's been a while since I've done maths, I have found that the answer is 3x4^n but I don't know how to get there

OpenStudy (anonymous):

\[4^{n+1}-4^{n}\]

OpenStudy (anonymous):

Wolfram Alpha says the answer is \[3 \times 4^{n}\]

ganeshie8 (ganeshie8):

hint : \(\large 4^{n+1} = 4^n \times 4^1\)

OpenStudy (anonymous):

I see absolutely no way of getting 3 from that o_0

OpenStudy (anonymous):

oh wait...

ganeshie8 (ganeshie8):

yes :) stare at it som more time :)

OpenStudy (anonymous):

Wouldn't it be 4 x 4^n then?

OpenStudy (anonymous):

4^1= 4

ganeshie8 (ganeshie8):

good going so far humm

OpenStudy (anonymous):

Why did wolfram alpha say the answer was 3 x 4^n? Which answer is right?

OpenStudy (anonymous):

Wolfram alpha output for reference: http://www.wolframalpha.com/input/?i=simplify+%284^%28n%2B1%29%29-%284^n%29

OpenStudy (anonymous):

oooops, I forgot the second part

ganeshie8 (ganeshie8):

\(4^{n+1} = 4^n \times 4^1 = 4\times 4^n\)

ganeshie8 (ganeshie8):

yes, subtract the second part

OpenStudy (isaiah.feynman):